A medical engineering company creates x-ray machines. The machines the company sold in 1995 were expected to last six years before breaking. To test how long the machines actually lasted, the company took a simple random sample of six machines. The company got the following results (in years) for how long the x-ray machines lasted: 8, 6, 7, 9, 5, and 7. Assume the distribution of the longevity of x-ray machines is normally distributed. Construct a 98% confidence interval for the average longevity of x-ray machines.
Answer #1

A medical engineering company creates x-ray machines. The machines the company sold in 1995 were expected to last six years before breaking.

The given data is

8, 6, 7, 9, 5, 7

Mean, ñ = $x = 42-7 g= ::Sample mean, 122=7] standard deviation :- sh(81-7) = 3 (–702 -12 = 14142 .:. Sample standard deviatilevel, The critical of t at significance X= 0.02 with df = 5 is t212 , df = 10.0212,5 - † 3.365 Margin g=error :- ME= tx 121d

Therefore the 98% confidence interval for the average longevity of x-ray machines is (5.06, 8.94)

Standard error = 0.58


answered by: ANURANJAN SARSAM